OpenStudy (anonymous):

PLEASE HELP!!! Rewrite each trig. expression in simpler form using trigonometric identities. a) 2sin5pcos5p b) sin pi/6 cos pi/4 - cos pi/6 sin pi/4

zepdrix (zepdrix):

Recall your Sine Double Angle Identity: \(\large\rm 2\sin(\color{orangered}{\theta})\cos(\color{orangered}{\theta})=\sin(2\color{orangered}{\theta})\) So if you have, \(\large\rm 2\sin(\color{orangered}{5p})\cos(\color{orangered}{5p})\) what will you end up with? Apply the identity.

OpenStudy (anonymous):

Thanks for helping by the way! So, I got sin (2 * 5p) = 2sin5pcos5p

zepdrix (zepdrix):

Good good good. simplifying to sin(10p), ya? :)

OpenStudy (anonymous):

Mhm! And then I divided sin5p from both sides.

OpenStudy (anonymous):

So I got sin2p = 2cos5p

zepdrix (zepdrix):

What :U no no, sin(10p) is your solution. Stop there.

zepdrix (zepdrix):

no no no no don't ever do that again XD

zepdrix (zepdrix):

ever :P

OpenStudy (anonymous):

Oh! That makes a lot of sense. Thank you so much!

zepdrix (zepdrix):

sin(10p)/sin(5p) does not equal sin(2p). The numbers are locked inside the sine operation. You can't divide like that :)

OpenStudy (anonymous):

So, I'd just keep whatever is on the left side?

zepdrix (zepdrix):

If you wrote the sin(10p) on the left side, then yes :) we usually draw our equals signs on the right side, but thats ok hehe

zepdrix (zepdrix):

For the other problem, apply your Angle Difference Formula for Sine, \(\large\rm \sin A\cos B-\sin B\cos A=\sin(A-B)\)

OpenStudy (anonymous):

So, then A would equal pi/6 and B would equal pi/4

OpenStudy (anonymous):

Would that just simplify to sin (pi/6-pi/4) ?

zepdrix (zepdrix):

Yes, but we would like to simplify the angle if we can :) We'll need a common denominator.

zepdrix (zepdrix):

Unless your teacher doesn't care, I dunno :D

OpenStudy (anonymous):

The common denominator would be 12. Then, it's 2pi/12 - 3pi/12, and you'd get -1pi/12?

zepdrix (zepdrix):

Mmm k great! =sin(-pi/12) Since sine is an `odd function`, the negative can pass outward like this, =-sin(pi/12) But that last step isn't too important :) don't worry about it.

OpenStudy (anonymous):

Alright! Thanks! Also, if you have the time, would you mind helping me to understand what an odd function is?

zepdrix (zepdrix):

It might be easier to first understand what an `even function` is. You're familiar with a parabola?

OpenStudy (anonymous):

Yes!

zepdrix (zepdrix):

Even functions are symmetric about the y-axis. Parabola is a good example, is has the same shape on the left and right side of the y-axis. But what's more important is that the negative always gets "squared" away in an even function. So even functions follow this property: \(\large\rm f(-x)=f(x)\) Example: If you plug x=-3 into the function, it should give you the same result as plugging x=3 into the function. \(\large\rm g(x)=(x)^2\) At x=2 and x-2, \(\large\rm g(-2)=(-2)^2=4\) \(\large\rm g(~~~2)=(~~~2)^2=4\)

OpenStudy (anonymous):

Oh! I see. So, an odd function would be like.. a cubic function?

zepdrix (zepdrix):

Yes, good good. \(\large\rm f(x)=(x)^3\) And you recall what happens when you multiply three negatives together?

OpenStudy (anonymous):

Yeah! It turns into a negative.

zepdrix (zepdrix):

\(\large\rm f(-2)=(-2)^3=-(2)^3=-f(2)\) Good good good. So we can take the negative out front like that.

zepdrix (zepdrix):

So odd function follows this property: \(\large\rm f(-x)=-f(x)\) the negative can come out in front of the function like that.

OpenStudy (anonymous):

And for the even functions, it wouldn't matter if the x value was positive or negative, it would still turn out positive?

zepdrix (zepdrix):

Yes, good :) Positive. And it turns out that Cosine is an example of an `even function`. So we could say that \(\large\rm \cos(-x)=\cos(x)\) The negative can go away. \(\large\rm \cos(-\pi/3)=\cos(\pi/3)\)

zepdrix (zepdrix):

While Tangent and Sine are `odd functions`.

OpenStudy (anonymous):

Would that be because cosine represents the x value?

OpenStudy (anonymous):

(regarding unit circle coordinates)

zepdrix (zepdrix):

Yes yes. I'm not sure if I can draw it for you since I'm on my laptop right now. Hmm

OpenStudy (anonymous):

That's alright, I understand

zepdrix (zepdrix):

So ya, they both have the same x-coordinate :)